OpenStudy (anonymous):

please help!!! WILL MEDAL FOR ANSWER!!! In October, Greg and Thomas had the same amount of money in their savings accounts. In November, Greg deposited $120 into his account. Thomas increased the money in his account by 20%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in October?

OpenStudy (anonymous):

will medal!!!

OpenStudy (anonymous):

please??

OpenStudy (studygurl14):

Set up two equations x= initial amount y = total amount in november Greg deposited 120 into account (after initial) y = x + 20 Thomas increased by 20%. y = x + (20%)x You can combine these equations into this x + 20 = x + 0.2x (since 20% = 0.2) Does that help?

OpenStudy (studygurl14):

@danburns123 Does the help you?

OpenStudy (anonymous):

yes

OpenStudy (studygurl14):

:)

OpenStudy (anonymous):

thanks!

OpenStudy (studygurl14):

anytime :)